OpenStudy (anonymous):

If f(t)=(5t-7/t)^4/7,find prime of f(x) using chain rule

OpenStudy (lgbasallote):

use power rule first

OpenStudy (anonymous):

How do I do that?

OpenStudy (lgbasallote):

\[\huge \frac{d}{dx} x^n \implies nx^{n-1}\]

OpenStudy (anonymous):

ok. So it would turn to (4/7)(5t-7/t)^-3/7

OpenStudy (lgbasallote):

wait yeah sorry

OpenStudy (lgbasallote):

now take the derivative of 5t - 7

OpenStudy (anonymous):

That would be just 5 right

OpenStudy (lgbasallote):

wait... it's 5t - 7/t

OpenStudy (lgbasallote):

so it's not just 5

OpenStudy (anonymous):

5-7?

OpenStudy (lgbasallote):

no.. derivative of 7/t is not 7

OpenStudy (lgbasallote):

\[\huge \frac 7t \implies 7t^{-1}\] use power rule on that

OpenStudy (anonymous):

Oh yeah because it come up so you have to make it in to a negative

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

so we are at (4/7) (5-7t^-1)^-3/7 do we multiply the (4/7)?

OpenStudy (lgbasallote):

you still have to derive 5t - 7/t...

OpenStudy (anonymous):

but wouldn't that be the second prime of f(x) ?

OpenStudy (lgbasallote):

it's chain rule

OpenStudy (lgbasallote):

\[\Large \frac{d}{dx} f(x)^n \implies n\cdot f(x) ^{n-1} f'(x)\]

OpenStudy (lgbasallote):

if you want to go algebra

OpenStudy (anonymous):

(4/7)(5-7t^-1)(5)

OpenStudy (lgbasallote):

we've been over this.. the derivative of 5t - 7/t is not 5

OpenStudy (lgbasallote):

how about i show you how to do the chain rule....it seems you're confused about it...

OpenStudy (lgbasallote):

\[\huge \frac d{da} (a^2 + 3a - 4)^3\] \[\huge \implies 3(a^2 + 3a - 4)^{3-1} \cdot \frac d{da} (a^2 + 3a - 4)\] \[\huge \implies 3(a^2 + 3a - 4)^2 \cdot (2a + 3)\] \[\huge \implies (6a + 9)(a^2 + 3a - 4)^2\] got it now?